OpenStudy (anonymous):

solve these linear equations -6x - 4y = 12 -2x + 5y = 4

OpenStudy (anonymous):

To solve such equations we make any variable common in both the equations and then subtract either of the equation. here, multiply second equation with 3 3 x -2x + 5y = 4 x 3 -6x + 15 y =12

OpenStudy (anonymous):

Now subtract this equation from first o -6x - 4y = 12 -(-6x + 15 y =12)

OpenStudy (anonymous):

now we get 0 - 19y = 0 or y = 0

OpenStudy (anonymous):

i guess now you can find the value of x by putting y = 0 in any one of the eqaution